OpenStudy (anonymous):

find three consecutive odd integers such that the product of the first and the second exceeds the third by 8.

OpenStudy (anonymous):

numbers are n,n+2, n+ 4 where n is odd and n(n+2) = n+4 + 8 n(n+2) = n + 12 solve this for n can you do that?

OpenStudy (anonymous):

n^2 + n - 12 = 0

OpenStudy (anonymous):

this equation has 2 roots , one even and one odd its the odd one u require

OpenStudy (anonymous):

do you need help solving this equation?

OpenStudy (anonymous):

lol yea im lost

OpenStudy (anonymous):

you need 2 numbers which when multiplied give -12 and when added give 1 these are 4 and -3 so we get (n+4)(n-3) = 0 so n = -4 or 3 n is given as odd so it must be 3 so your 3 consecutive odd numbers are 3, 5 and 7

OpenStudy (anonymous):

oooooh thanx:)

OpenStudy (anonymous):

yw